### v1

# ViM-VQ: Efficient Post-Training Vector Quantization for Visual Mamba

Abstract
Visual Mamba networks (ViMs) extend the selective space state model (Mamba) to various vision tasks and demonstrate significant potential. Vector quantization (VQ), on the other hand, decomposes network weights into codebooks and assignments, significantly reducing memory usage and computational latency to enable ViMs deployment on edge devices. Although existing VQ methods have achieved extremely low-bit quantization (e.g., 3-bit, 2-bit, and 1-bit) in convolutional neural networks and Transformer-based networks, directly applying these methods to ViMs results in unsatisfactory accuracy. We identify several key challenges: 1) The weights of Mamba-based blocks in ViMs contain numerous outliers, significantly amplifying quantization errors. 2) When applied to ViMs, the latest VQ methods suffer from excessive memory consumption, lengthy calibration procedures, and suboptimal performance in the search for optimal codewords. In this paper, we propose ViM-VQ, an efficient post-training vector quantization method tailored for ViMs. ViM-VQ consists of two innovative components: 1) a fast convex combination optimization algorithm that efficiently updates both the convex combinations and the convex hulls to search for optimal codewords, and 2) an incremental vector quantization strategy that incrementally confirms optimal codewords to mitigate truncation errors. Experimental results demonstrate that ViM-VQ achieves state-of-the-art performance in low-bit quantization across various visual tasks.
1 Introduction
Recent advancements have sparked growing research interest in state space models (SSMs). [19]. Modern SSMs are particularly adept at capturing long-range dependencies and benefit from parallel training capabilities. Several SSM-based methods, including the structured state space sequence model (S4) [12] and S4D [13], have been introduced to efficiently handle long sequence data across various tasks, owing to their convolutional computations and near-linear complexity. The recent development of Mamba [11] extends the SSM framework by incorporating time-varying parameters and proposing a hardware-aware algorithm for highly efficient training and inference. This has spurred increasing interest in adapting Mamba to vision tasks [44, 25, 17, 41], giving rise to the concept of Visual Mamba networks (ViMs). Among these ViMs, Vision Mamba (Vim) and VMamba have gained particular attention due to their effectiveness and unique architectures.
Model quantization compresses deep neural networks (DNNs), enabling their deployment and inference on resource-constrained devices with limited computational resources. Quantization methods are generally classified into uniform quantization (UQ) [20, 9, 29, 22, 39, 10, 23, 32] and vector quantization (VQ)  [33, 28, 1, 42, 35, 4, 8, 34]. UQ reduces the precision of network weights, while VQ clusters network weights into codebooks and corresponding assignments. Meanwhile, post-training quantization (PTQ) [27, 31] is the most commonly used technique, as it rapidly calibrates the quantization parameters using a small calibration set, without the need for time-consuming fine-tuning. Both UQ and VQ have been widely applied to convolutional neural networks [30] and Transformer-based networks [37], with research on UQ in ViMs already underway [40, 2]. Nevertheless, the feasibility of achieving extreme compression via VQ for ViMs remains unexplored.
Figure 1: Visualization of the weight distribution and quantization error for the Mamba-based blocks in advanced Visual Mamba network. (a) VMamba [25] comprises numerous Mamba-based [11] blocks (e.g., VSS blocks). (b) The weights of these blocks exhibit significant outliers. (c-d) Although conventional vector quantization (VQ) outperforms uniform quantization (UQ) at low-bit precision, substantial quantization errors still severely degrade network performance. (e) Our ViM-VQ enables the quantized network to achieve performance comparable to that of the original network.
We first identify several critical challenges encountered when applying vector quantization (VQ) to ViMs: 1) Outliers in ViMs’ Weights. As shown in Figure 1, the weights of the Mamba-based [11] blocks in ViMs exhibit significant outliers. The presence of significant outliers leads to considerable quantization errors, resulting in unsatisfactory accuracy. 2) Limitations of Existing VQ Methods. As depicted in Figure 2, existing VQ methods for visual networks suffer from substantial calibration steps, significant GPU memory consumption, and suboptimal performance when applied to multi-layered networks such as ViMs. 3) Truncation Error from Assignment Conversion. Figure 3 illustrates that weight average methods suffer from truncation errors introduced during the soft-to-hard assignment conversion, causing degradation of inference performance relative to calibration.
To overcome these challenges, we propose ViM-VQ, an efficient post-training vector quantization method for Visual Mamba networks (ViMs). ViM-VQ introduces a fast convex combination optimization algorithm that efficiently initializes and updates both the convex combinations and the convex hulls, rapidly searching for optimal codewords while significantly reducing computational and memory overhead. Additionally, ViM-VQ adopts an incremental vector quantization strategy to incrementally confirm optimal codewords to mitigate truncation errors. Our major contributions are summarized as follows:
  • Our ViM-VQ efficiently quantizes ViMs and achieves superior performance while significantly reducing both calibration time and GPU memory requirements.
  • To the best of our knowledge, our ViM-VQ is the first vector quantization framework specifically designed for Visual Mamba networks (ViMs). We analyze ViMs’ weight distribution and address the unique challenges of applying VQ to ViMs through targeted improvements.
  • Compared to advanced VQ method, experiments demonstrate that our ViM-VQ achieves better performance in low-bit quantization across various visual tasks.
Figure 2: Comparison of different vector quantization (VQ) algorithms. (a) Conventional VQ algorithms [14, 33, 28] assign the optimal codeword to each weight sub-vector based solely on distance, making them unsuitable for the task. (b) Global weighted average VQ algorithms [1, 42] apply soft assignments to all codewords and leverage differentiable weighted averaging to identify optimal codewords based on the objective function. However, these methods require substantial calibration efforts and memory overhead, particularly in multi-layered structures such as Visual Mamba networks (ViMs) [44, 25]. (c) Local weighted average VQ algorithms [4] mitigate issues of global approaches by applying soft assignments to only a few nearest codewords, yet their performance remains suboptimal due to the limited scope. (d) Our ViM-VQ overcomes the above challenges and adopts a fast convex combination optimization algorithm to efficiently initialize and optimize both the convex combinations and the convex hulls to search for optimal codewords. Furthermore, ViM-VQ employs an incremental vector quantization strategy to incrementally confirm optimal codeword and compensate for truncation errors.
2 Related Work
Visual Mamba Networks.
Mamba [11] is an advanced structured state space model that significantly enhances the capabilities of state space models (SSMs) for processing sequential data. It transforms the parameters of the structured state space model (S4) [12] into learnable functions, introducing a parallel scanning method to improve efficiency. By overcoming the local perception limitations of convolutional neural networks (CNNs) [30] and the quadratic computational complexity of Transformers [37], Visual Mamba networks (ViMs) [24] have found widespread success across a range of visual tasks. Vision Mamba (Vim) [44] represents the first use of Mamba in computer vision, utilizing bidirectional SSMs for global feature modeling and position embeddings to enhance spatial awareness. Subsequently, VMamba [25] introduces a cross-scan module to address direction-sensitive challenges. LocalMamba [17] improves performance by incorporating local inductive biases, while PlainMamba [41] adopts a non-hierarchical structure for better multi-scale integration. Mamba-ND [21] extends the Mamba architecture to multi-modal data by simply altering the sequence order. Despite the impressive performance of ViMs in computer vision tasks, their large size remains a significant obstacle for deployment on edge devices with limited resources.
Model Quantization.
Model quantization is an effective technique to compress deep neural networks by reducing their model size and improving inference efficiency. Current quantization methods can be broadly classified into two main categories [35]: uniform quantization (UQ), which converts model weight from high to low precision, and vector quantization (VQ), which decomposes model weight into a codebook and corresponding assignments. Most quantization approaches are implemented via post-training quantization (PTQ) [27, 31], which requires only a small calibration dataset. In the UQ domain, GPTQ [10] proposes a layer-wise quantization method that leverages approximate second-order information to quantize weights with minimal accuracy loss. Recent studies have begun exploring UQ methods specifically tailored for ViMs. PTQ4VM [2] introduces per-token static quantization and joint learning of smoothing scale and step size to mitigate outliers. MambaQuant [40] applies the Hadamard transformation to standardize the variance across channels. In the VQ domain, PQF [28] utilizes rate-distortion theory to reorder weights, thereby reducing clustering errors. DKM [1] integrates a differentiable K-Means formulation into the objective function to preserve the network’s accuracy. JLCM [42] builds upon DKM by adjusting the gradient computation for assignments to avoid incorrect updates. VQ4DiT [4] identifies that fixed initial assignments can lead to suboptimal updates, proposing an improved search algorithm to dynamically select more appropriate codewords. Although these methods have shown promising results in CNNs and Transformers, their performance under extremely low-bit quantization settings remains limited in ViMs. To the best of our knowledge, our approach is the first post-training vector quantization solution explicitly designed for ViMs.
3 Preliminaries
3.1 Mamba
Mamba [11] is a state space model (SSM) inspired by continuous system, mapping a 1-D sequence $x(t)\in\mathbb{R}\mapsto y(t)\in\mathbb{R}$ through a hidden state $h(t)\in\mathbb{R}^{\mathtt{N}}$. This system uses $\mathbf{E}\in\mathbb{R}^{\mathtt{N}\times\mathtt{N}}$ as the evolution matrix and $\mathbf{B}\in\mathbb{R}^{\mathtt{N}\times 1}$, $\mathbf{P}\in\mathbb{R}^{1\times\mathtt{N}}$ as the projection matrices. The continuous system works as follows: $h^{\prime}(t)=\mathbf{E}h(t)+\mathbf{B}x(t)$ and $y(t)=\mathbf{P}h(t)$. Mamba introduces a timescale parameter $\mathbf{\Delta}$ to transform $\mathbf{E}$ and $\mathbf{B}$ to discrete matrices $\mathbf{\overline{E}}$ and $\mathbf{\overline{B}}$. This transformation employs the zero-order hold method, defined as:
$\mathbf{\overline{E}}$ $=\exp(\mathbf{\Delta}\mathbf{E}),$ (1)
$\mathbf{\overline{B}}$ $=(\mathbf{\Delta}\mathbf{E})^{-1}(\exp(\mathbf{\Delta}\mathbf{E})-\mathbf{I})\cdot\mathbf{\Delta}\mathbf{B}.$
With these discrete matrices, the system dynamics and the output computed via a global convolution can be respectively rewritten as:
$y_{t}$ $=\mathbf{P}h_{t}=\mathbf{P}(\mathbf{\overline{E}}h_{t-1}+\mathbf{\overline{B}}x_{t}),$ (2)
$\mathbf{y}$ $=\mathbf{x}*\mathbf{\overline{K}}=\mathbf{x}*(\mathbf{P}\mathbf{\overline{B}},\mathbf{P}\mathbf{\overline{E}}\mathbf{\overline{B}},\dots,\mathbf{P}\mathbf{\overline{E}}^{\mathtt{M}-1}\mathbf{\overline{B}}),$ (3)
where $\mathtt{M}$ is the length of the input sequence $\mathbf{x}$, and $\overline{\mathbf{K}}\in\mathbb{R}^{\mathtt{M}}$ is a structured convolutional kernel.
3.2 Visual Mamba Networks
Vision Mamba (Vim) [44] and VMamba [25] are the most commonly used Visual Mamba networks (ViMs) [24]. Inspired by ViT [7] and BERT [5], Vim is the first network to extend Mamba to vision tasks. Vim incorporates position embeddings for annotating image sequences and utilizes bidirectional state space models to efficiently compress visual representations. VMamba achieves superior performance compared to Vim. The architectural overview of the tiny-scale VMamba is illustrated in Figure 1(a). The input image is first partitioned into patches by a stem module, yielding an initial 2D feature map. Without additional positional embeddings, VMamba progressively constructs hierarchical features across multiple stages, each comprising a downsampling layer and several Visual State Space (VSS) blocks analogous to Mamba-based blocks.
3.3 Vector Quantization
Vector quantization (VQ) is a more flexible approach than uniform quantization (UQ), especially suitable for extremely low-bit quantization scenarios. VQ typically employs clustering algorithms such as K-Means [15] to decompose the weight matrix $\mathbf{W}\in\mathbb{R}^{o\times i}$ into codebook and assignments. Specifically, $\mathbf{W}$ is divided into $d$-dimensional sub-vectors $\mathbf{W}=\{w_{o,i/d}\}$, where $o\times i/d$ is the total number of sub-vectors. These sub-vectors are then represented using a codebook $\mathbf{C}\in\mathbb{R}^{k\times d}$, where $c(k)$ refers to the $k$-th codeword of length $d$. The assignments $\mathbf{A}=\{a_{o,i/d}\}$ indicate the index of the codeword that best represents each sub-vector. Finally, the quantized weight $\widehat{\mathbf{W}}$ is reconstructed by replacing each $w_{o,i/d}$ with $c(a_{o,i/d})$:
$\widehat{\mathbf{W}}=\mathbf{C}[\mathbf{A}]=\begin{bmatrix}c(a_{1,1})&c(a_{1,2})&\cdots&c(a_{1,i/d})\\ c(a_{2,1})&c(a_{2,2})&\cdots&c(a_{2,i/d})\\ \vdots&\vdots&\ddots&\vdots\\ c(a_{o,1})&c(a_{o,2})&\cdots&c(a_{o,i/d})\end{bmatrix}.$ (4)
Assignments $\mathbf{A}$ can be stored using $\frac{o\times i}{d}\times\log_{2}k$ bits, and codebook $\mathbf{C}$ can be stored using $k\times d\times 32$ bits (typically negligible). Therefore, the quantization bit for quantizing 32-bit origin weights can be computed as $\frac{d\times 32}{\log_{2}k}$.
4 Method
4.1 Motivation
Quantizing the weights of Visual Mamba networks (ViMs) to extremely low-bit precision significantly reduces memory usage and accelerates inference, making ViMs suitable for deployment on resource-constrained edge devices. However, as illustrated in Figure 1, ViMs’ weights contain numerous outliers, leading to substantial quantization errors when applying conventional uniform quantization (UQ) or vector quantization (VQ), ultimately leading to model performance degradation. The relatively smaller quantization error observed in VQ indicates its superiority over UQ in extremely low-bit quantization scenarios [35, 4], which motivates our exploration of applying VQ to ViMs. As illustrated in Figure 2, existing VQ approaches exhibit the following limitations when applied to ViMs:
Conventional VQ. Deep compression [14], ATB [33], and PQF [28] fail to adequately adjust the distribution of quantized weights to support extremely low-bit quantization. Consequently, these methods have primarily been effective for smaller networks (e.g., ResNet-18 [16]) and show limited scalability for larger models like ViMs.
Global Weighted Average VQ. DKM [1] and JLCM [42] require substantial GPU memory for storing and updating soft assignment ratios and incur significant calibration steps. These limitations render them impractical for multi-layered architectures such as ViMs. For instance, as indicated in Table 5, using a $256\times 4$ codebook for 2-bit quantization significantly increases the number of learnable parameters by $4256\times o\times i/$ 4 per layer, potentially causing out-of-memory issues on resource-constrained devices.
Local Weighted Average VQ. Although VQ4DiT [4] addresses the memory and computational overhead issues present in earlier methods, its performance remains suboptimal due to the limited scope of candidate codewords (see Figure 3. Moreover, initializing soft assignments with uniform values results in a substantial initial discrepancy between the quantized weights and the original weights, causing gradient oscillations during optimization.
Truncation Error from Assignment Conversion. Figure 3 and Table 5 demonstrate that existing weighted average VQ methods neglect the truncation errors introduced during the conversion from soft to hard assignments when determining the optimal codewords. This oversight leads to noticeable inference accuracy degradation compared to calibration results.
Figure 3: Distributions of the largest soft assignment ratios after applying existing weighted average vector quantization (using 4 soft assignments per sub-vector and a codebook of size $256\times 8$). The largest ratios not close to 1.0 demonstrate two issues: (1) a few optimal codewords lie outside current candidate codewords, leading to suboptimal performance, and (2) directly confirming optimal codewords based on largest ratios leads to significant truncation errors.
Therefore, we propose ViM-VQ, a fast post-training vector quantization method specifically tailored for extremely quantizing ViMs, which consists of two innovative components: 1) To overcome the challenges of high GPU memory usage, multiple calibration steps, and suboptimal performance encountered by weighted average VQ methods, we develop a fast convex combination optimization algorithm. This algorithm initializes convex combinations of several nearby candidate codewords and soft assignment ratios. Both the convex combinations and the convex hulls are rapidly optimized, significantly accelerating the search for optimal codewords. 2) To eliminate the truncation error resulting from assignment conversion, we develop an incremental vector quantization strategy, in which soft assignments with high ratios are transitioned from soft to hard to confirm optimal codewords in each calibration iteration. This strategy ensures that parts still undergoing calibration can compensate for truncation errors, preventing these errors from affecting the entire network.
Figure 4: Incremental vector quantization strategy of ViM-VQ. Soft assignments with high ratios (e.g., $\geq$ 0.99) undergo a soft-to-hard conversion to confirm the optimal codewords, while other components that are still being calibrated compensate for truncation errors.
4.2 Fast Convex Combination Optimization
[…]
$\mathcal{C}_{o,i/d}=\{c_{o,i/d}^{1},\dots,c_{o,i/d}^{n}\}=\arg\min_{k}^{n}\|w_{o,i/d}-c(k)\|_{2}^{2},$ (5)
where $\mathcal{C}_{o,i/d}$ is candidate codewords of sub-vector $w_{o,i/d}$. Next, we compute soft assignment ratios $\mathcal{R}{o,i/d}$ for these candidate codewords:
$\mathcal{R}_{o,i/d}=\{r_{o,i/d}^{1},\dots,r_{o,i/d}^{n}\}=\{\dots,\frac{\exp\left({z_{o,i/d}^{n}}\right)}{\sum_{j=1}^{n}\exp\left({z_{o,i/d}^{j}}\right)}\},$ (6)
subject to the softmax constraint $\sum_{n}r_{o,i/d}^{n}=1$. Here, $z_{o,i/d}^{n}$ represents the unnormalized score for each ratio. In this way, the learnable $\mathcal{C}_{o,i/d}$ along with the learnable $\mathcal{R}_{o,i/d}$ form differentiable convex combinations that produces the quantized weight $\widehat{\mathbf{W}}$ as follows:
$\widehat{\mathbf{W}}\!=\!\\ \begin{bmatrix}\widehat{w}_{1,1}\!=\!\sum\mathcal{C}_{1,1}\mathcal{R}_{1,1}&\!\cdots\!&\widehat{w}_{1,i/d}\!=\!\sum\mathcal{C}_{1,i/d}\mathcal{R}_{1,i/d}\\ \widehat{w}_{2,1}\!=\!\sum\mathcal{C}_{2,1}\mathcal{R}_{2,1}&\!\cdots\!&\widehat{w}_{2,i/d}\!=\!\sum\mathcal{C}_{2,i/d}\mathcal{R}_{2,i/d}\\ \vdots&\ddots&\vdots\\ \widehat{w}_{o,1}\!=\!\sum\mathcal{C}_{o,1}\mathcal{R}_{o,1}&\!\cdots\!&\widehat{w}_{o,i/d}\!=\!\sum\mathcal{C}_{o,i/d}\mathcal{R}_{o,i/d}\\ \end{bmatrix}.$ (7)
For the initialization of each $\mathcal{R}_{o,i/d}$, we optimize $z_{o,i/d}^{n}$ via gradient descent to minimize the reconstruction error between $\widehat{\mathbf{W}}$ and $\mathbf{W}$:
$z_{o,i/d}^{n}=z_{o,i/d}^{n}-\eta\frac{\partial\mathcal{L_{\text{init}}}}{\partial z_{o,i/d}^{n}},\quad\mathcal{L_{\text{init}}}=\|\mathbf{W}-\widehat{\mathbf{W}}\|_{2}^{2},$ (8)
where $\eta$ denotes the learning rate. This initialization ensures that $\widehat{\mathbf{W}}$ is initially close to $\mathbf{W}$, thus preventing severe gradient oscillations during subsequent calibration.
Therefore, the convex combination optimization facilitates the quantized sub-vectors’ search for the optimal codeword. How can we ensure that optimal codewords are contained within the convex hulls? The quantized sub-vectors are often distant from those irrelevant candidate codewords with near-zero ratios, which indicates that the optimal codeword is not located near these irrelevant candidates. Consequently, we can replace them with new candidate codewords that are closest to the quantized sub-vector:
$c_{o,i/d}^{n}|_{r_{o,i/d}^{n}\approx 0}=\arg\min_{k}\|\widehat{w}_{o,i/d}-c(k)\|_{2}^{2}\notin\mathcal{C}_{o,i/d}$ (9)
where $r_{o,i/d}^{n}\approx 0$ is smaller than $\lambda$ (e.g., 1e-2). This adaptive replacement procedure dynamically shifts the convex hulls towards more optimal codewords without introducing significant perturbations to the network’s outputs.
The key differences between the fast convex combination optimization algorithm and previous works can be summarized as follows:
  • Unlike global weighted average methods [1, 42], which incur high calibration costs and memory overhead, or local weighted average methods [4], which yield suboptimal quantization, our approach efficiently balances these aspects. Our algorithm utilizes fewer codewords and soft assignment ratios to reduce GPU memory usage and calibration steps. Simultaneous updates of convex combinations and convex hulls accelerate the search for optimal codewords, achieving superior performance.
  • Rather than using hard assignments and approximated gradients [14, 36, 18, 33, 28], our algorithm leverages differentiable convex combinations directly based on the objective function of the task.
  • Our algorithm is a differentiable process and does not require additional computation such as Hessian trace [6, 3, 26, 8] or SVD [18, 38, 45].
4.3 Incremental Vector Quantization
As shown in Figure 3, weighted average VQ algorithms enforce a one-time soft-to-hard transition in all assignments to confirm the optimal codewords, which introduces truncation errors into quantized weights:
$\sum\left|c_{o,i/d}^{n}|_{r_{o,i/d}^{n}\approx 1}-\sum\mathcal{C}_{o,i/d}\mathcal{R}_{o,i/d}\right|>0.$ (10)
Table 5 further demonstrates that these truncation errors cause deviations in intermediate features within each block of ViMs, ultimately accumulating and leading to decreased accuracy in the final output.
Inspired by INQ [43], we propose an incremental vector quantization strategy. As illustrated in Figure 4, our strategy confirms candidate codewords with soft assignment ratios close to 1.0 (e.g. 0.99) as optimal codewords at each calibration step:
$\widehat{w}_{o,i/d}=c_{o,i/d}^{n}|_{r_{o,i/d}^{n}\approx 1}.$ (11)
Compared to previous methods that perform a one-time assignment conversion, our incremental strategy significantly reduces truncation errors at each calibration step. This allows the weights that remain in calibration to effectively recover the network’s accuracy. Moreover, our method ensures consistent network performance between the calibration and inference phases.
The incremental vector quantization strategy introduces three objective functions to efficiently identify and confirm optimal codewords. Specifically, given the input $\mathbf{x}$ and the target $\mathbf{y}$ from the calibration dataset for the quantized ViMs $\mathbf{\epsilon}_{q}$, the task objective function is defined as:
$\mathcal{L}_{t}=\mathbb{E}_{\mathbf{x},\mathbf{y}}\left[\|\mathbf{y}-\mathbf{\epsilon}_{q}(\mathbf{x})\|_{2}^{2}\right].$ (12)
To ensure consistency in the Mamba-based block outputs between the original and quantized networks, we employ a block-wise knowledge distillation (BKD) objective:
$\mathcal{L}_{bkd}=\mathbb{E}_{\mathbf{x}}\left[\sum_{l}\|b_{fp}^{l}(\mathbf{x})-b_{q}^{l}(\mathbf{x})\|_{2}^{2}\right],$ (13)
where $b_{fp}^{l}$ and $b_{q}^{l}$ represent the features of $l$-th block from $\mathbf{\epsilon}_{fp}$ and $\mathbf{\epsilon}_{q}$, respectively.
Additionally, we introduce a regularization term $\mathcal{L}_{r}$ to encourage rapid convergence of the differentiable convex combinations:
$\mathcal{L}_{r}=\left(\frac{d}{o\times i}\right)\times\sum_{o,i/d}\left(\mathcal{R}_{o,i/d}\times\left(1-\mathcal{R}_{o,i/d}\right)\right).$ (14)
The final combined objective function $\mathcal{L}$ is formulated as:
$\mathcal{L}=\mathcal{L}_{t}+\mathcal{L}_{bkd}+\begin{cases}\mathcal{L}_{r},&\text{if }\mathcal{L}_{r}^{(t)}>\mathcal{L}_{r}^{(t-1)},\\[5.0pt] 0,&\text{otherwise},\end{cases}$ (15)
where $\mathcal{L}_{r}$ is selectively included only when it increases compared to the previous iteration. This selective incorporation ensures stable updates and convergence of the differentiable convex combinations, ultimately accelerating the search for optimal codewords.
5 Experiment
Our experiments are conducted on two Visual Mamba networks (ViMs) [24]: Vision Mamba (Vim) [44] and VMamba [25]. The experiments cover four main computer vision tasks: image classification, semantic segmentation, object detection, and instance segmentation. Our primary goal is to investigate the trade-off between model weight quantization and accuracy. Therefore, activation quantization is disabled in all experiments. We sample 100 images from each training category to form the calibration set, using a batch size of 128. All experiments are conducted on a single NVIDIA A6000 GPU with 48GB of memory.
Hyperparameters: The codebook sizes for each layer in the 3/2/1-bit quantization are set to $2^{6}\times 2$, $2^{8}\times 4$, and $2^{8}\times 8$, respectively. Soft assignments ratios are updated by the Adamax optimizer with a fixed learning rate of $5\times 10^{-2}$ without additional hyperparameter tuning. Dataset processing and other hyperparameters are identical to those of the original ViMs. Consistent with all baselines, we quantize only the linear projection layers within the ViMs blocks, as these layers contribute significantly to memory usage.
Baselines: We compare our method against a broad range of uniform quantization (UQ) and vector quantization (VQ) methods. UQ methods include basic uniform quantization (Round-to-Nearest), GPTQ [10], and recent ViMs-specific quantization techniques such as Mambaquant [40] and PTQ4VM [2]. The recent VQ methods designed for visual networks include the basic vector quantization (K-Means [15]), PQF [28], DKM [1], and VQ4DiT [4].
Model Method Top-1 Accuracy (%)
FP 3-bit 2-bit 1-bit
Vim-T $\text{UQ}^{\dagger}$ 76.07 52.84 0.16 -
GPTQ 56.74 0.17 -
Mambaquant 57.21 0.18 -
PTQ4VM 57.29 0.18 -
$\text{VQ}^{\dagger}$ 67.56 33.18 0.22
PQF 71.71 61.84 4.20
DKM 73.64 71.25 68.29
VQ4DiT 73.57 71.04 67.72
ViM-VQ (Ours) 74.79 72.17 69.93
Vim-S $\text{UQ}^{\dagger}$ 80.48 72.31 0.16 -
GPTQ 74.45 0.20 -
Mambaquant 74.57 0.24 -
PTQ4VM 74.68 0.25 -
$\text{VQ}^{\dagger}$ 77.59 65.96 0.37
PQF 78.46 73.28 10.93
DKM 79.70 77.96 70.38
VQ4DiT 79.37 77.87 69.73
ViM-VQ (Ours) 80.10 78.66 72.02
Vim-B $\text{UQ}^{\dagger}$ 81.88 76.17 2.06 -
GPTQ 76.81 2.47 -
Mambaquant 77.01 3.49 -
PTQ4VM 77.17 3.62 -
$\text{VQ}^{\dagger}$ 77.78 70.88 5.208
PQF 78.76 76.02 27.40
DKM 80.02 79.18 74.57
VQ4DiT 79.73 78.86 74.37
ViM-VQ (Ours) 80.34 79.46 75.58
Table 1: Quantization results of Vision Mamba (Vim) for image classification on ImageNet-1K. $\text{UQ}^{\dagger}$ and $\text{VQ}^{\dagger}$ represent basic uniform quantization and vector quantization, respectively.
5.1 Image Classification
Table 1 presents the image classification performance on three different-scale Vision Mamba (Vim-T/S/B) evaluated on ImageNet-1K. Uniform quantization (UQ) methods collapse in 2-bit quantization due to significant quantization errors, even with the latest Vim-specific approaches, such as Mambaquant and PTQ4VM. Across all three networks and various bit-widths, our ViM-VQ consistently achieves state-of-the-art quantization performance, maintaining the highest Top-1 accuracy. Specifically, for ViM-T, ViM-VQ achieves accuracies of 74.79% (3-bit), 72.17% (2-bit), and 69.93% (1-bit), clearly outperforming other advanced VQ methods like DKM and VQ4DiT. Similar trends appear in ViM-S and ViM-B, where ViM-VQ maintains performance very close to the full-precision baseline, demonstrating robustness even under extremely low-bit quantization.
Bit Method
mIoU
mIoU
mAcc aAcc
FP Swin-T 44.5 45.8 - -
ConvNeXt-T 46.0 46.7 - -
VMamba-T 47.9 48.8 59.2 82.4
2 PQF 42.7 43.6 58.2 79.7
DKM 44.1 44.8 58.7 80.6
VQ4DiT 43.9 44.7 58.4 80.6
ViM-VQ (Ours) 44.4 45.4 58.9 81.0
1 PQF 28.5 29.0 39.9 72.6
DKM 37.2 37.6 52.8 76.6
VQ4DiT 36.2 36.8 51.9 76.2
ViM-VQ (Ours) 41.1 41.8 56.0 77.3
Table 2: Quantization results of tiny-scale VMamba (VMamba-T) for semantic segmentation on ADE20K. ‘SS’ and ‘MS’ denote single-scale and multi-scale testing, respectively. UperNet is utilized as the segmentation framework.
[…]
FP Swin-T 42.7 - 39.3 -
ConvNeXt-T 44.2 - 40.1 -
VMamba-T 47.3 69.3 42.7 66.4
2 PQF 44.6 66.5 40.6 63.7
DKM 45.7 68.2 41.5 64.8
VQ4DiT 45.5 67.9 41.4 64.7
ViM-VQ (Ours) 46.0 68.8 41.8 65.7
1 PQF 33.1 53.0 31.2 50.5
DKM 37.5 58.0 35.4 55.4
VQ4DiT 36.6 56.9 34.0 54.3
ViM-VQ (Ours) 40.9 60.7 35.8 57.9
[…]
5.2 Semantic Segmentation
Table 2 presents the semantic segmentation results of VMamba-T on the ADE20K dataset using UperNet as the segmentation framework. At 2-bit quantization, ViM-VQ achieves an mIoU of 44.4% (single-scale) and 45.4% (multi-scale), attaining performance comparable to full-precision Swin-T and ConvNeXt-T. Even under aggressive 1-bit quantization, ViM-VQ maintains an mIoU above 40%, demonstrating its robustness in extremely low-bit settings.
5.3 Object Detection and Instance Segmentation
Table 3 presents the object detection and instance segmentation performance on MSCOCO using Mask R-CNN as the detection framework. At 2-bit quantization, ViM-VQ consistently outperforms full-precision Swin-T and ConvNeXt-T, achieving 46.0% box AP and 41.8% mask AP. Even under the extreme condition of 1-bit quantization, ViM-VQ significantly mitigates performance degradation, obtaining 40.9% box AP and 35.8% mask AP.
Bit Method Top-1 Acc (%)
FP Vim-T 76.07
2 Baseline ($\text{VQ}^{\dagger}$) 33.18
+ 1) Combination Optimization 71.42 (38.24$\uparrow$)
+ 2) Incremental Quantization (Ours) 72.17 (0.75$\uparrow$)
Table 4: Ablation results for ViM-VQ for 2-bit quantization of tiny-scale Vision Mamba (Vim-T) on ImageNet-1K. $\text{VQ}^{\dagger}$ represents basic vector quantization.
$\mathbf{C}$ Method $n$ $E$ Mem Acc-C Acc-I
256 $\times$ 4 DKM 256 14 25 GB 71.98 71.25
VQ4DiT 4 2 11 GB 71.33 71.04
ViM-VQ 4 2 11 GB 72.17 72.17
ViM-VQ 16 4 12 GB 72.31 72.31
ViM-VQ 64 8 14 GB 72.34 72.34
Table 5: Calibration efficiency of 2-bit quantization of tiny-scale Vision Mamba (Vim-T) on ImageNet-1K. $\mathbf{C}$, $n$, $E$, ”Mem” represent the codebook size, the number of candidate codewords, the number of epochs, and the GPU memory used for calibration, respectively. ”Acc-C” and ”Acc-I” represent the Top-1 accuracy (%) during calibration and inference.
Bit Method Resolution Size (MB) CUDA Time
FP Vim-B 224$\times$224 1452 n/a 336s
2 ViM-VQ 224$\times$224 92 $\times$ 357s
224$\times$224 92 $\checkmark$ 328s
Table 6: Inference time of 2-bit base-scale Vision Mamba (Vim-B) on ImageNet-1K validation set on a single NVIDIA A6000 GPU. ’Size’ represents the model size. ’CUDA’ denotes whether the CUDA kernel for matrix–vector multiplication is used.
5.4 Ablation Study
Table 4 presents an ablation study on the contribution of each component in our proposed ViM-VQ framework. The basic VQ method suffers from substantial quantization errors due to weight outliers, leading to a significant accuracy drop to 33.18%. By introducing the differential convex combination optimization algorithm, the accuracy improves by 38.24%, surpassing the performance of both DKM and VQ4DiT. This result highlights that updating both the convex combinations and the convex hulls effectively facilitates the search for the optimal codewords. Furthermore, incorporating the incremental vector quantization strategy provides an additional 0.75% improvement, demonstrating that the truncation error induced during the confirmation of optimal codewords is effectively mitigated.
As shown in Table 5, we report the calibration efficiency for 2-bit quantization of Vim-T on ImageNet-1K. DKM requires the most epochs and GPU memory during calibration, which can easily lead to out-of-memory issues. Although VQ4DiT significantly reduces computational overhead, its performance remains suboptimal due to the limited scope of candidate codewords. Moreover, DKM and VQ4DiT neglect the truncation error, resulting in a noticeable decrease in inference accuracy compared to calibration. In contrast, ViM-VQ outperforms DKM and VQ4DiT while maintaining performance consistency, indicating that it strikes a better balance between memory usage, calibration time, and final accuracy.
We develop a customized CUDA kernel for efficient matrix–vector multiplication. In our implementation, the codebooks are stored in constant GPU memory, while the assignments are fully loaded into CUDA blocks, enabling each thread to decode its corresponding assignment in parallel. The input is sequentially multiplied by the decoded weights, with the results accumulated in registers. As shown in Table 6, our CUDA kernel reduces inference time from 357s to 328s, bringing the speed of the quantized model similar to that of the original ViMs.
6 Conclusion
In this work, we analyze key challenges encountered when applying vector quantization (VQ) to Visual Mamba networks (ViMs). First, significant weight outliers in ViMs lead to severe quantization errors. Second, existing VQ methods applied to ViMs require substantial calibration steps, high memory consumption, and yield suboptimal performance. Additionally, truncation errors further degrade inference performance. We propose an efficient post-training VQ method tailored for ViMs, termed ViM-VQ. ViM-VQ incorporates a fast convex combination optimization algorithm for optimal codeword search and an incremental vector quantization strategy to confirm optimal codewords while mitigating truncation errors. Experimental results demonstrate that ViM-VQ significantly outperforms existing approaches across diverse visual tasks, offering a practical solution for deploying ViMs on edge devices.